# Random Alloy Codes and the Fundamental Limits of Coded Distributed Tensors

### v6

@@ Abstract
Tensors are a fundamental operation in distributed and are commonly distributed into multiple parallel tasks for large datasets. Stragglers and other failures can severely impact the overall completion time. Recent works in coded computing provide a novel strategy to mitigate stragglers with coded tasks, with an objective of minimizing the number of tasks needed to recover the overall result, known as the recovery threshold. However, we demonstrate that this strict combinatorial definition does not directly optimize the probability of failure.
@@ I Introduction
Machine learning has become a dominant tool in the broader computing community. Its success has been due to the availability of large datasets and more recently the use of hardware optimized to perform multi-linear functions such as the GPU and the TPU, which are a fundamental building block for a large number of deep neural network architectures. If the dataset has many data points then the overall computation, or job, is distributed as tasks amongst workers, which model a distributed network of computing devices. This solution creates a new problem; stragglers and other faults can severely impact the performance and overall training time.
@@ I-B Background
If $A$ is a $x\times z$ matrix with entries in $\mathbb{F}^{P\times R}$ (equivalently, $A$ is a $xP\times zR$ matrix that has been partitioned into $x\times z$ parts) and $B$ is a $z\times y$ matrix with entries in $\mathbb{F}^{R\times Q}$, then, if $T$ is the matrix multiplication operator takes the form $T:(\mathbb{F}^{P\times R})^{x\times z}\times(\mathbb{F}^{R\times Q})^{z\times y}\rightarrow(\mathbb{F}^{P\times Q})^{x\times y}$, then we define the tensor rank as the smallest number $r$ of $x\times z$-dimensional linear functionals $E^{i}_{0}:(\mathbb{F}^{P\times R})^{x\times z}\rightarrow\mathbb{F}^{P\times R}$, $z\times y$-dimensional linear functionals $E^{i}_{1}:(\mathbb{F}^{R\times Q})^{z\times y}\rightarrow\mathbb{F}^{R\times Q}$, and coefficients $D_{i}\in\mathbb{F}$ such that $T(A,B)=\sum_{i\in[r]}D_{i}E^{i}_{0}(A)E^{i}_{1}(B).$ This modification is a straightforward modification of the usual definition and is implicitly used for recursive fast matrix multiplication algorithms. In practice, the parameters $P,Q,R$ are determined by the amount of memory/computing power available at the worker nodes. Suppose we are given some family of subsets, $\mathcal{R}\subset[n]$, where $n$ is the number of workers, which we take to be the recoverable subsets. Then we have a family of coefficients $D_{i}^{\mathcal{R}}$ so that the coded matrix multiplication tensor is given by
@@
The works [4, 5, 6] established the use of coding theory for distributed coded matrix-vector and vector-vector multiplication; shortly after and concurrently, the works [1, 7] further improved on these techniques by using alignment to come up with efficient constructions for general matrix-matrix multiplication. In particular, the authors of [1] established the equivalence of the recovery threshold for general partitions to the tensor rank of matrix multiplication (which they call the bilinear complexity). The most similar approaches to our work are the codes presented in [3] and the codes in [1] as well as the impossibility results in [8]. The works [6, 9, 10, 7, 11, 12, 13, 14, 15] also worked on improving coded distributed matrix multiplication. Further work has been extended to include batch matrix multiplication as well [16, 17, 18, 19]. The work [20] considered the extra constraint of private/secure multiplication and the work [21] considered a trade-off between the size of the entries of the matrix and the recovery threshold. The works [22, 23] considered coding for distributed gradient descent in machine learning applications.
@@ II Code Constructions and Analysis
if we use standard matrix notation. If $d_{1}+...+d_{\ell}\geq d_{1}\cdot...\cdot d_{\ell}$, then we have that the probability that $G$ is full rank is equal to
$\prod_{i\in[d]}(1-q^{i-d_{1}-...-d_{\ell}}).$ (4)
@@
$\mathcal{N}(k,d)=\prod_{i\in[k]}(q^{d_{1}+...+d_{\ell}}-q^{i})$
by induction. For the base case, a $1\times d$ matrix $g$ is full rank iff its first and only row $(g)^{1}$ is a non zero vector. There are a total of
$\mathcal{N}(1,d)=q^{d_{1}+...+d_{\ell}}-1$
non-zero rank 1 vectors since $(g)^{1}_{i_{1},...,i_{\ell}}=(g_{1})^{1}_{i_{1}}\cdot...\cdot(g_{\ell})^{1}_{i_{\ell}}$ for some $g_{i}\in\mathbb{F}^{d_{i}}$ by definition and there are $q^{d_{1}}$ choices for $(g_{1})^{1}$, $q^{d_{2}}$ choices for $(g_{2})^{1}$, and so on and we subtract out the 0 vector.
@@
$\mathcal{N}(k,d)=\prod_{i\in[k]}(q^{d_{1}+...+d_{\ell}}-q^{i}).$
To make a full rank $k+1$ row matrix we must first choose $k$ many linearly independent rows of flattened rank 1 matrices and then for the choice of the $(k+1)^{\mathrm{th}}$ row we must subtract out all possible linear combinations of the previous $k$ rows. There are $(q^{d_{1}+...+d_{\ell}}-q^{k+1})$ many such linear combinations and thus an induction gives us that
$\mathcal{N}(k+1,d)=\mathcal{N}(k,d)(q^{d_{1}+...+d_{\ell}}-q^{k+1})\\ =\prod_{i\in[k+1]}(q^{d_{1}+...+d_{\ell}}-q^{i}).$
@@
$\frac{\mathcal{N}(k,d)}{\mathcal{T}(k,d)}=\prod_{i\in[k]}\frac{(q^{d_{1}+...+d_{\ell}}-q^{i})}{q^{d_{1}+...+d_{\ell}}}=\prod_{i\in[d]}(1-q^{i-d_{1}-...-d_{\ell}}),$
@@ II-C An Impossibility Theorem
There do not exist linear outer product codes (of neither the deterministic nor random type) if $d_{1}+...+d_{\ell}<d_{1}\cdot...\cdot d_{\ell}$.
@@
Recall in the proof of Theorem 1 that we proved that the total number of full rank $(k+1)\times d$ matrices that can be created by flattening outer products satisfies the recursion
$\mathcal{N}(k+1,d)=\mathcal{N}(k,d)(q^{d_{1}+...+d_{\ell}}-q^{k+1}),$
which counts the number of $k+1$ linearly independent sets of flattened rank 1 matrices by taking a set of $k$ many such matrices and subtracting out all the possible linear combinations of the set. However, plugging in $d$ gives us $\mathcal{N}(d,d)=0,$ and thus there cannot exist $d$ many linearly independent rank 1 tensors if $d_{1}+...+d_{\ell}>d$, as was needed to be shown. ∎
Remark 1.
Theorem 4 is surprising since there are many works [5, 7, 20] that perform order 2 outer products as one of their main contributions and claim to work over all fields. But we have see from the statement that already if $A$ is a $3\times 1$ matrix and $B$ is a $2\times 1$ matrix that there can be no such linear code over such a finite field.
@@ II-D Locally Random Alloy Codes
is sufficient. Algebraic manipulation and the facts that $\exp\left(-\frac{\left(\frac{\lambda}{1+\lambda}-p\right)^{2}\chi}{2p\left(1-p\right)}\right)<1$ and $c<1\implies 1-c\geq\log(1-c)$ together imply that

### v7

@@ Abstract
Tensors are a fundamental operation in distributed computing, e.g., machine learning, that are commonly distributed into multiple parallel tasks for large datasets. Stragglers and other failures can severely impact the overall completion time. Recent works in coded computing provide a novel strategy to mitigate stragglers with coded tasks, with an objective of minimizing the number of tasks needed to recover the overall result, known as the recovery threshold. However, we demonstrate that this strict combinatorial definition does not directly optimize the probability of failure.
@@ I Introduction
Machine learning has become a dominant tool in the broader computing community. Its success has been due to the availability of large datasets and more recently the use of hardware optimized to perform multi-linear functions such as GPU’s and TPU’s, which are a fundamental building block for a large number of deep neural network architectures. If the dataset has many data points then the overall computation, or job, is distributed as tasks amongst workers, which model a distributed network of computing devices. This solution creates a new problem; stragglers and other faults can severely impact the performance and overall training time.
@@ I-B Background
If $A$ is a $x\times z$ matrix with entries in $\mathbb{F}^{P\times R}$ (equivalently, $A$ is a $xP\times zR$ matrix that has been partitioned into $x\times z$ parts) and $B$ is a $z\times y$ matrix with entries in $\mathbb{F}^{R\times Q}$, then, if $T$ is the matrix multiplication operator takes the form $T:(\mathbb{F}^{P\times R})^{x\times z}\times(\mathbb{F}^{R\times Q})^{z\times y}\rightarrow(\mathbb{F}^{P\times Q})^{x\times y}$, then we define the tensor rank as the smallest number $r$ of $x\times z$-dimensional linear functionals $E^{i}_{0}:(\mathbb{F}^{P\times R})^{x\times z}\rightarrow\mathbb{F}^{P\times R}$, $z\times y$-dimensional linear functionals $E^{i}_{1}:(\mathbb{F}^{R\times Q})^{z\times y}\rightarrow\mathbb{F}^{R\times Q}$, and coefficients $D_{i}\in\mathbb{F}$ such that $T(A,B)=\sum_{i\in[r]}D_{i}E^{i}_{0}(A)E^{i}_{1}(B).$ This modification is a straightforward modification of the usual definition and is implicitly used for recursive fast matrix multiplication algorithms. In practice, the parameters $P,Q,R$ are determined by the amount of memory/computing power available at the worker nodes.
Suppose we are given some family of subsets, $\mathcal{R}\subset[n]$, where $n$ is the number of workers, which we take to be the recoverable subsets. Then we have a family of coefficients $D_{i}^{\mathcal{R}}$ so that the coded matrix multiplication tensor is given by
@@
The works [4, 5, 6] established the use of coding theory for distributed coded matrix-vector and vector-vector multiplication; shortly after and concurrently, the works [1, 7] further improved on these techniques by using alignment to come up with efficient constructions for general matrix-matrix multiplication. In particular, the authors of [1] established the equivalence of the recovery threshold for general partitions to the tensor rank of matrix multiplication (which they call the bilinear complexity). The most similar approaches to our work are the codes presented in [3] and the codes in [1] as well as the impossibility results in [8]. The works [6, 9, 10, 7, 11, 12, 13, 14, 15] also worked on improving coded distributed matrix multiplication. Further work has been extended to include batch matrix multiplication as well [16, 17, 18, 19]. The work [20] considered the extra constraint of private/secure multiplication and the work [21] considered a trade-off between the size of the entries of the matrix and the recovery threshold, and parallel integer multiplication [22]. The works [23, 24] considered coding for distributed gradient descent in machine learning applications.
@@ II Code Constructions and Analysis
if we use standard matrix notation. If $d_{1}+...+d_{\ell}\geq d_{1}\cdot...\cdot d_{\ell}$, then we have that the probability that $G$ is full rank is greater than or equal to
$\prod_{i\in[d]}\left[\left(\prod_{j\in[\ell]}1-q^{-d_{j}}\right)-q^{i-d_{1}-...-d_{\ell}}\right].$ (4)
@@
$\mathcal{N}(k,d)\geq\prod_{i\in[k]}\left[\left(\prod_{j\in[\ell]}q^{d_{j}}-1\right)-q^{i}\right]$
by induction. For the base case, a $1\times d$ matrix $g$ is full rank iff its first and only row $(g)^{1}$ is a non zero vector. There are a total of $\mathcal{N}(1,d)=\prod_{j\in[\ell]}(q^{d_{j}}-1)$ non-zero rank 1 vectors since $(g)^{1}_{i_{1},...,i_{\ell}}=(g_{1})^{1}_{i_{1}}\cdot...\cdot(g_{\ell})^{1}_{i_{\ell}}$ for some $g_{i}\in\mathbb{F}^{d_{i}}$ by definition and there are $q^{d_{1}}-1$ choices for $(g_{1})^{1}$, $q^{d_{2}}-1$ choices for $(g_{2})^{1}-1$, and so on and we subtract out the 0 vector ($\ell$ times).
@@
$\mathcal{N}(k,d)\geq\prod_{i\in[k]}\left[\left(\prod_{j\in[\ell]}q^{d_{j}}-1\right)-q^{i}\right].$
To make a full rank $k+1$ row matrix we must first choose $k$ many linearly independent rows of flattened rank 1 matrices and then for the choice of the $(k+1)^{\mathrm{th}}$ row we must subtract out all possible linear combinations of the previous $k$ rows; this is bounded by $q^{k+1}$. There are $(\prod_{j\in[\ell]}q^{d_{j}}-1)-q^{k+1}$ many such linear combinations and thus an induction gives us that
$\mathcal{N}(k+1,d)\geq\mathcal{N}(k,d)(\prod_{j\in[\ell]}q^{d_{j}}-1)-q^{k+1})\\ \geq\prod_{i\in[k+1]}\left[\left(\prod_{j\in[\ell]}q^{d_{j}}-1\right)-q^{i}\right].$
@@
$\frac{\mathcal{N}(k,d)}{\mathcal{T}(k,d)}=\prod_{i\in[k]}\frac{\left(\prod_{j\in[\ell]}q^{d_{j}}-1\right)-q^{i}}{q^{d_{1}+...+d_{\ell}}}\\ =\prod_{i\in[k]}\left[\left(\prod_{j\in[\ell]}1-q^{-d_{j}}\right)-q^{i-d_{1}-...-d_{\ell}}\right],$
@@ II-C An Impossibility Theorem
There do not exist linear outer product codes (of neither the deterministic nor random type) if $\left(\prod_{j\in[\ell]}1-q^{-d_{j}}\right)\leq q^{1-d_{1}-...-d_{\ell}}\cdot d_{1}\cdot...\cdot d_{\ell}$.
@@
A modification of the proof of Theorem 1 gives us that the total number of full rank $(k+1)\times d$ matrices that can be created by flattening outer products satisfies the recursion
$\mathcal{N}(k+1,d)\leq\mathcal{N}(k,d)\left(\left(\prod_{j\in[\ell]}q^{d_{j}}-1\right)-q(k+1)\right),$
which counts the number of $k+1$ linearly independent sets of flattened rank 1 matrices by taking a set of $k$ many such matrices and subtracting out all the possible linearly dependent choices. However, plugging in $d$ gives us $\mathcal{N}(d,d)/\mathcal{T}(d,d)\leq 0,$ and thus there cannot exist $d$ many linearly independent rank 1 tensors if $\left(\prod_{j\in[\ell]}1-q^{-d_{j}}\right)<q^{1-d_{1}-...-d_{\ell}}\cdot d_{1}\cdot...\cdot d_{\ell}$. ∎
@@ II-D Locally Random Alloy Codes
is sufficient. Algebraic manipulation and the facts that
$\exp\left(-\frac{\left(\frac{\lambda}{1+\lambda}-p\right)^{2}\chi}{2p\left(1-p\right)}\right)<1$
and $c<1\implies 1-c\geq\log(1-c)$ together imply that